GDP at market prices R397bn
Net primary income payments to the rest of
the world
R37bn
Indirect taxes R23bn
Subsidies R11bn
Consumption of fixed capital R32bn
1 Calculate the value of gross national income (GNI) at market prices.
2 Calculate the value of net national product (NNI) at market prices.
3. Calculate the value of net national income (NNI) at factor cost.
1
Expert's answer
2020-09-28T10:42:27-0400

a) Gross National Income (GNI) = GDP + Net payments to the world + Taxes - Subsidies - Consumption of fixed assets


= R (397 "+" 37 "+" 23 "-" 11 "-" 32) bn


= R 414bn


b) Net National Product(NNI) at Market Prices.

=GDP fc − Consumption of fixed capital − Net Primary income payments to the rest of the world.


=[(397"+" 32) "-" 32"-" 37]


=R 360bn


c) NNI at Factor cost =NNI at Market Price +Subsidies "-" Indirect taxes


=(360+11"-"23 )


=R 348bn